how do u fin the equivelant to the exponent like 2 to the power of 20 over 2 to the power of 4
Answer by stanbon(75887)

how do u fin the equivelant to the exponent like 2 to the power of 20 over 2 to the power of 4
--------------
When you divide you subtract exponents, as follows:
2^20/2^4 = 2^(20-4)= 2^16
Cheers,
Stan H.